# UNB StepUP: A footStep database for gait analysis and recognition using Underfoot Pressure

### v2

@@ Methods
Gait Feature Extraction and Validation
The present dataset’s technical quality and reliability were evaluated by analyzing different common gait representations and parameters. This evaluation was conducted in comparison to other publicly available datasets that utilize pressure or force measurements for gait, with findings detailed in the Technical Validation section. Within the literature on plantar pressure, it is typical to employ a range of two-dimensional spatial representations to distill specific characteristics of each pixel’s values over time, such as its peak pressure, pressure-time integral, contact duration, and time-to-maximum [21]. This approach synthesizes all pertinent details from the time series of pressure images throughout a stance phase into a comprehensive image. Among these features, peak pressure images are probably the most widely employed in the literature, and are computed from each footstep’s 3D tensor as the maximum pressure experienced by each pixel (i.e., sensor) during the stance. Peak pressure images were used here to illustrate the variations in shod pressure patterns (Fig. 8) as well as the differences in sensor density across datasets (Fig. 13). Biomechanical analyses also frequently involve time series signals due to the temporal nature of gait patterns, such as vertical ground reaction force (GRF) profiles, and center of pressure (COP) trajectories. GRF time series, representing the forces exerted by the ground on the foot throughout the stance phase, are computed as the sum of pressures across all pixels at each time step. COP time series are defined in the mediolateral (ML, $x$-axis) and anteroposterior (AP, $y$-axis) directions, and are calculated as the pressure-weighted average of the foot’s coordinates at each time step during the stance. These two time series were used for quality assessment and comparison with other public datasets (Fig. 14). Lastly, several spatiotemporal gait parameters were derived from the pressure recordings, including gait speed, cadence, step length, step width, and toe-out angle. With the exception of toe-out angle, these features were extracted using the distances between 3D bounding box coordinates ($time$, $x$, $y$) from consecutive footsteps. Toe-out angle was computed as the angle of the first principal component axis of the footstep (i.e., along the longest dimension of the foot) with respect to the long axis of the tile grid, and is negative for inward rotations and positive for outward rotations. These spatiotemporal parameters were compared to those from the CASIA-D dataset (Figs. 15-16).
Data Records
The official documentation for the dataset and all required scripts can be found on its GitHub page: https://github.com/UNB-StepUP/StepUP-P150. The StepUP-P150 dataset is accessible for download on figshare[22] (https://doi.org/10.6084/m9.figshare.28143686) and contains the raw recordings, two variants of extracted and preprocessed footstep data, along with detailed metadata for each individual footstep. The files are organized as shown in Fig. 11. Located at the top level, the spreadsheet named ‘participant_metadata.csv’ contains demographic information, anthropometric measurements, types of footwear, and other possible influences on the gait patterns of each for the 150 participants. More comprehensive information regarding the metadata fields is available in Table 4. Footstep pressure data is also structured into separate folders at the top level, with each corresponding to a single participant. Each folder is named using the pattern XXX, with ‘XXX’ representing a unique, randomly assigned ID that encompasses all participants (001-150). Within each participant’s folder, four (second-level) sub-folders represent the various footwear conditions: ‘BF’, ‘ST’, ‘P1’, and ‘P2’. Inside these, there are seven further (third-level) sub-subfolders designated for distinct standing balance and walking experiment trials: ‘S1’, ‘S2’, ‘S3’, ‘W1’, ‘W2’, ‘W3’, or ‘W4’ (refer to Fig. 4 and the Experimental Protocol section for more information).
For each of the three standing balance trials (S1, S2, and S3), there are two corresponding files. The ‘trial.{npz,mat}’ file houses the raw, unprocessed 30-second recording from the trial, represented as a 3D tensor with dimensions of nominally $3000\times 720\times 240$ (frames, height, width; note that the number of frames may vary slightly). The processed version of the recording is available as ‘preprocessed.{npz,mat}’, representing a 3D tensor with dimensions of $3000\times 180\times 180$ (frames, height, width). These files are available in both NumPy (.npz) and MATLAB (.mat) file formats, each of which uses a dictionary-like structure for variable storage, with the tensors accessible under the top-level key arr_0. Unlike the balance trials, each folder for the walking trials (W1, W2, W3, and W4) includes four files. The raw, unprocessed 90-second pressure data is available in ‘trial.{npz,mat}’, containing a 3D tensor with dimensions of $9000\times 720\times 240$ (frames, height, width; again, the number of frames may vary). The footsteps were extracted using two distinct preprocessing methods and are provided as ‘pipeline_1.{npz,mat}’ and ‘pipeline_2.{npz,mat}’, which are elaborated upon in the Footstep Normalization section. Each tensor is 4D with dimensions $n_{footsteps}\times 101\times 75\times 40$ (samples, frames, height, width), where $n_{footsteps}$ specifies the number of footsteps in a given trial. These tensors are all accessible via the key arr_0. Each individual footstep’s metadata is provided in a file named ‘metadata.csv’, containing details such as the spatiotemporal location in the initial recording, classification labels (e.g., left or right foot, walking direction, outlier), and various parameters obtained during preprocessing (e.g., rotation angle, foot length and width, mean GRF). Table 5 lists the metadata fields and their explanations.
Technical Validation
Evaluation of the present database, along with other publicly available pressure-based gait databases (Table 1), is conducted in this section using the 5 V’s of big data as a framework: volume, variety, velocity, veracity, and value [23]. Although pressure-based floor sensor gait datasets have not yet reached the large scale of big data seen in other gait modalities such as video- or wearable-based approaches, these data characteristics are employed to evaluate the current state of this field, offering insight for the advancement of future databases and datasets in this specialized area.
Volume
The SFootBD database contains approximately 20,000 footsteps total from 127 individuals [10], making it the largest database of its kind until the introduction of the current database, UNB StepUP. The newly introduced database starts with an initial dataset named UNB StepUP-P150, described in this paper, which includes over 200,000 footsteps collected from 150 individuals. This makes it more than ten times larger than the SFootBD in terms of total footsteps (Table 1). The SFootBD database contains three benchmark datasets named B1, B2, and B3 which simulate three unique data-driven security environments, each with a varying number of users and training footsteps per user: airport security screenings (limited training data), workplace environments (moderate training data), and home settings (extensive training data). Considering the nature of the SFootBD database, where some individuals provided multiple footstep recordings during several sessions, while others contributed only a few, the benchmarks were determined by the number of participants that provided a specific number of footstep samples. Consequently, they proposed simulating three different scenarios: an airport security checkpoint with 40 users and 80 training footsteps per user, a workplace with 15 users and 400 training footsteps each, and a home environment with 5 users and 1000 training footsteps each, with testing restricted to 10 footsteps per user across all scenarios. By adopting a similar structure for access checkpoints and providing 80 training steps for each user, the UNB StepUP-P150 dataset can increase its number of users from 40 to 150. On the other hand, for home security systems involving five users, the UNB StepUP-P150 dataset can elevate the training process for each user to include approximately 1,400 steps. In fact, the UNB StepUP-P150 dataset can emulate various scenarios involving up to 150 users, each contributing 1,400 steps on average for model training and evaluation.
In addition to the number of subjects and steps, the UNB StepUP database features a substantially larger floor-sensing platform and an enhanced sensing resolution (or greater sensor density) compared to other publicly available databases. The 120 cm $\times$ 360 cm sensing tile grid used for the StepUP-P150 database is the largest yet, which allows the collection of natural walking behaviour over a span of 4-6 consecutive footsteps, in contrast to the 1-4 steps recorded by other databases (Fig. 12). Additionally, the StepUP-150 database employed pressure-sensing tiles featuring a spatial resolution of 5 mm $\times$ 5 mm, offering the highest sensor density. This represents a more than 50% increase along the axis of walking compared to the next two highest-resolution underfoot pressure databases: CAD WALK and CASIA-D (Fig. 13). A much higher sensing resolution translates to a significantly larger number of voxels per footstep. Until now, high-resolution underfoot pressure data was only available in much smaller datasets, such as the CASIA-D dataset, which is arguably the most well-known in this field and contains just over 3,000 footsteps from 118 individuals [9].
Variety
Different factors, known as gait covariates, can influence gait patterns. These include walking speed, footwear, elapsed time, carried load, walking environment, clothing, direction of gait and foot angle with respect to the sensors, and dual-task gait. In previous research on underfoot pressures, the main challenges often stem from variations in walking speed and the types of footwear worn. Specifically, the CASIA-D dataset consists of two distinct sub-datasets, with each one concentrating on a different covariate: one on walking speed (normal and fast speeds) and the other on footwear types (barefoot and two kinds of shoes, including running shoes, Chinese cloth shoes, and leather shoes). While these datasets allow for the exploration of each factor’s individual impact, in real-world scenarios, these factors frequently interact with each other. In contrast, the SFootBD database aimed to replicate real-world scenarios by allowing participants to walk at their natural pace and wear any type of footwear, though annotated labels were not provided for changes in footwear or other conditions. Although this database facilitates the investigation of the effects of combined factors, it does not allow the evaluation of individual covariates. This limitation also hinders the development of specialized algorithms tailored for applications with variable covariates, for instance, access control in areas where individuals inherently remove their shoes, such as some religious places or specific airport security checkpoints, where the primary concern is barefoot samples. The UNB StepUP-P150 experiment simultaneously examined the two factors by incorporating four different walking speeds (normal, fast, slow, and slow-to-stop, simulating the approach to a turnstile at a controlled access point) along with four types of footwear conditions (barefoot, standard shoes, and two personal pairs), resulting in a total of 16 unique covariate conditions.
Additionally, ongoing experiments that will extend the UNB StepUP database consist of two additional studies. Building on the first study that led to the UNB StepUP-P150 dataset, the second study adds two more factors: time elapsed and carried load. Meanwhile, the third study includes four additional factors: walking setting, worn apparel, direction of gait and foot angle with respect to the sensors, and dual-task gait, which integrates cognitive activities. Individuals who took part in the first study were invited to participate in the subsequent second and third studies. This methodology allows for the integration of temporal effects (across different days) into these studies, while preserving specific conditions associated with the two main factors: walking speed and type of footwear. As a result, the data collection protocols led to 33 distinct covariate conditions for the second study and 24 for the third. At the time of submission, footstep data have been collected from 40 participants for the second study and 21 participants for the third. This data collection process is ongoing, and these datasets will be made publicly available as new components of the UNB StepUP database in the future, where they may be referred to as the UNB StepUP-CV dataset. The designation CV refers to the gait covariate problem.
Velocity
Several important points concerning the speed of data collection and processing should be highlighted. Firstly, a 90-second walk can lead to anywhere between 50 and 120 steps, averaging approximately 90 steps, depending on the walking speed. This allows biometric enrollment to be completed quickly, taking as little as one to two minutes. With the current protocol involving 16 unique conditions and related tasks like measuring body dimensions, changing footwear, and incorporating breaks between trials, gathering around 1,400 footsteps is completed in under 90 minutes. Secondly, the segmentation and annotation of individual footsteps can be achieved through methods suitable for real-time execution. In the UNB StepUP-P150 dataset, the SORT algorithm [16], which is designed for real-time tracking, is instrumental in the segmentation of each individual footstep. In addition, automatic algorithms were employed to identify left and right footfalls, incomplete or outlier footprints, and to distinguish between standing and walking footsteps. Thirdly, the UNB StepUP-P150 already features footstep data for 150 individuals, but efforts are continuing to further increase the number of participants. These additional subjects may also serve an important role in representing unknown users and impostors in biometric research, possibly as part of a future competition at an international conference. In addition, data collection for the second and third studies is ongoing. To our knowledge, the collection of the publicly available pressure-based gait datasets listed in Table 1 have been completed, with no ongoing data collection.
As part of the project that enabled the UNB StepUP database, an additional instrumented runway with a $4\times 6$ grid of pressure-sensing tiles was installed at the secure entry point of the Cyber Center Building in Fredericton, New Brunswick, Canada. This setup in a leading cybersecurity center in Canada has facilitated the collection of footstep data in a genuine high-security setting, reflecting real-world usage patterns such as arriving for work, taking lunch breaks, and leaving. The system has been active for the past 3 years, recording data from more than 130,000 footsteps collected from more than 90 consented individuals. This real-world dataset, the first known such dataset of its kind, incorporates a mix of uncontrolled variables and two common traffic routes onto the tiles and is planned for future public release as part of the UNB StepUP database, where it may be referred to as the UNB StepUP-RW dataset. The designation RW refers to a non-controlled real-world setting, similar to “in-the-Wild” datasets used for vision-based gait recognition.
Veracity
There are several noteworthy aspects to highlight. Firstly, the UNB StepUP-P150 dataset includes different versions of pre-processing pipelines for footstep pressure data, offering multiple options for users. Scripts for these preprocessing pipelines, which combine three components— spatial normalization, temporal normalization, and amplitude normalization (Table 3)—are also provided. The GRF and COP time series derived from the preprocessed StepUP-P150 dataset exhibited an average waveform comparable to those found in other gait datasets that use force or pressure data, such as the combined GaitRec and Gutenberg dataset [6, 7], the dataset produced by Derlatka and Partieniuk [24], and CASIA-D (Fig. 14). Secondly, concerning labels, although automated algorithms have facilitated annotation and allowed real-time application in practical implementation, a rigorous manual inspection was performed on the UNB StepUP-P150 dataset. This was done to ensure that all labels were accurately corrected, involving two rounds of review with a minimum of two examiners, as detailed in the Data Quality Assessment section. Video recordings frequently served as an extra resource to aid in verifying potential label inconsistencies (presumed mistakes that may occur during experimental procedures). This approach was used in the manual inspection of the current dataset and other public datasets such as CASIA-D and SFootBD. Thirdly, this dataset further incorporated two metadata tags, Incomplete and Outlier, enabling users to omit incomplete or low-quality footstep samples from their analysis. The Incomplete metadata is mainly used to filter out partial or incomplete footsteps, ensuring that poor-quality incomplete footsteps are removed while allowing low quality but complete footsteps to remain. On the other hand, the Outlier metadata focuses on eliminating low-quality footsteps in general, encompassing many incomplete ones (though not all) and complete footsteps that may fall below quality standards. These two sets of metadata identified some common footsteps for exclusion but also flagged others, allowing users the choice to exclude footsteps based on one set, the other, both, or their overlap (Fig. 10).
Value
The initial UNB StepUP-P150 dataset from the UNB StepUP database shows marked improvements compared to other publicly available pressure-based gait databases and datasets in many aspects. These include, among others, a larger data size, a wider range of confounding factors, ongoing data collection, and meticulous data annotation, as discussed above, all contributing to the improved quantity and quality of the data. The larger size of the current dataset and the extensive range of covariates provide greater variation and a more thorough representation of gait patterns across general populations, which can be seen in spatio-temporal gait parameters such as step length, step width, cadence, and toe-out angle (Fig. 15). As an illustration, the toe-out angle range in CASIA-D is considerably more restricted compared to that of StepUP-P150. The limitation arises because CASIA-D requires subjects to walk along a strict and narrow straight line. In contrast, in real-world scenarios, individuals usually have more freedom in choosing their walking path and direction, leading to increased variability in the “viewing angle” covariate, as evidenced in the UNB StepUp-P150 dataset (Fig. 15). An additional example is shown in Fig. 16. In this figure, the scatter plots illustrate the average step length and step width (or support base) for 13 participants from CASIA-D and 150 participants from UNB StepUP-P150, each wearing two distinct pairs of personal shoes. Using a person identification model developed earlier using the CASIA-D dataset [25], these two gait parameters stood out prominently and are included in the feature sets that allow the gait recognition model to reach an accuracy of 90.5% with a single step and 99.5% with five steps. This model is trained with one pair of the participant’s personal shoes and assessed with shoes that were absent during training, a second pair of the participant’s personal shoes, which represents one of the most challenging scenarios for pressure-based gait recognition. However, this result may no longer hold, as illustrated in Fig. 16. Although the 13 subjects in the CASIA-D dataset can be perfectly distinguished using only these two gait parameters, numerous individuals in the StepUP-P150 dataset possess overlapping values, making recognition more challenging. These examples suggest that the UNB StepUP-P150 dataset has the potential to establish a new, more rigorous benchmark for pressure-based gait recognition.
Although the UNB StepUP database focuses mainly on pressure-sensitive floor sensors, it will also incorporate other modalities, such as video data and digital footprints/shoeprints. Before starting the collection of footstep data, the left and right footprints of each participant were scanned along with two pairs of shoeprints from their personal shoes using a flatbed digital scanner. Seven cameras were strategically placed to capture views from the front, the sides, and across all four corners. The video footage was recorded synchronously with the footstep pressure data. Once available, these data could serve as a standalone method for biometric recognition or be used in combination with other modalities for multi-modal recognition. These datasets will be made publicly available in the future as part of the UNB StepUP database, where they may be differentiated with labels such as the UNB StepUP-V150 and UNB StepUP-DS150 datasets. The potential designations V150 and DS150 would indicate that the data are related to video recordings and digital scans, gathered from 150 individuals, respectively. For the CASIA-D dataset, video footage was gathered using a side camera, while the SFootBD datasets employed both a front-facing and a side camera, capturing either one or two viewing angles. By comparison, StepUP-V150 will include views from 7 different cameras.
The UNB StepUP database was initially created to support gait recognition in biometric systems, such as those used to identify and verify individual gait patterns. To address demographic imbalances within the dataset, the UNB StepUP-P150 dataset incorporates participants with varied age, gender, race/ethnicity, and body type, effectively representing the diverse spectrum commonly observed in human gait. The creation of the current gait dataset, while addressing demographic bias within biometrics, provides the UNB StepUP-P150 dataset with extensive research prospects in the fields of gait analysis and recognition, offering applications that surpass merely biometric purposes. Subsequent research using the UNB StepUP-P150 dataset could focus on one of the research questions outlined below:
  • Statistical models of normative walking gait by utilizing foot pressure patterns
  • Differences in pressure-based gait patterns between demographic subgroups, categorized by factors like gender and age differences
  • Differences in pressure-based gait patterns resulting from external factors such as different walking speeds and types of footwear
  • Novel machine learning and deep learning models for gait recognition across various classification challenges
  • Novel approaches for the segmentation, alignment, and/or registration of plantar pressure images
  • Benchmark study to evaluate state-of-the-art techniques for gait analysis and gait recognition
@@
First, footstep data, along with supplementary metadata, are saved in *.npz and *.mat files, which can be readily imported and utilized with standard Python and MATLAB toolboxes. Scripts are provided to facilitate straightforward data import for Python and MATLAB users. These scripts are available on the dataset’s GitHub page (https://github.com/UNB-StepUP/StepUP-P150). The complete collection of UNB StepUP-P150 files can be accessed on figshare[22] (https://doi.org/10.6084/m9.figshare.28143686). Refer to the Data Records section for additional information regarding folder and file names. Secondly, raw, unprocessed pressure data is provided for each trial, allowing users to perform any preferred downstream processing methods. The dataset is accompanied by the custom script used for data processing, detailed in the Footstep Normalization section. While the dataset comprises preprocessed data from two distinct pipelines for quick prototyping and analysis, the script offers the flexibility to experiment with numerous other combinations. The authors have shared this script on the dataset’s GitHub page, allowing further optimization of the script. Third, scripts are provided to extract multiple standard gait features (e.g., peak pressure images, COP and GRF time series, and spatiotemporal gait parameters like step length and width) presented in the Technical Validation section. These are supplied to assist in benchmarking and in integrating or comparing the current dataset with other datasets. Finally, as part of an ongoing project, at least four other datasets will be made publicly available as part of the UNB StepUP database, including:
  • UNB StepUP-V150: The seven-camera video recordings dataset was collected along with the current footstep pressure data, obtained from 150 participants in the first protocol.
  • UNB StepUP-DS150: The dataset consisting of digital footprints and shoeprints was gathered from 150 participants who participated in the first protocol.
  • UNB StepUP-CV: The gait covariate dataset encompasses an extensive list of covariate factors, such as walking speed, type of footwear, cross-day gait, load carriage, walking environment, clothing, viewing angle, and dual-task gait, drawn from the second and third protocols.
  • UNB StepUP-RW: The dataset of footsteps from real-world scenarios was gathered by employing identical pressure-sensitive tiles at the secure entrance of the Cyber Centre Building, a premier cybersecurity facility in Canada.
Together, these gait datasets will offer numerous possibilities for researchers specializing in gait analysis and recognition.
@@
Custom scripts designed for processing and technical validation are provided on the dataset’s companion GitHub page (https://github.com/UNB-StepUP/StepUP-P150) to support ongoing improvements and optimizations. It should be noted that these scripts were created using MATLAB (The MathWorks, Inc., Natick, Massachusetts, United States, 2023a) and Python (Python Software Foundation, 3.11). The Python scripts require specific libraries, which are listed in the requirements.txt file. This file enables library installation through The Python Package Index (PyPI, https://pypi.org) or the Anaconda software distribution (2024.02, https://www.anaconda.com). For detailed instructions regarding the use and execution of the custom code, please consult the GitHub repository’s README file.

### v3

@@ Methods
Comparison to Existing Datasets
The dataset’s technical quality and reliability were evaluated by analyzing different common gait representations and parameters. This evaluation was conducted in comparison to other publicly available datasets that use pressure or force measurements for gait.
Spatial Representations
Within the literature on plantar pressure, it is typical to employ a range of two-dimensional spatial representations to distill specific characteristics of each pixel’s values over time, such as its peak pressure, pressure-time integral, contact duration, and time-to-maximum [22]. This approach synthesizes all pertinent details from the time series of pressure images throughout a stance phase into a comprehensive image. Among these features, peak pressure images are probably the most widely employed in the literature, and are computed from each footstep’s 3D tensor as the maximum pressure experienced by each pixel (i.e., sensor) during the stance. Peak pressure images were used here to illustrate the variations in unshod and shod pressure patterns (Fig. 8) as well as the differences in sensor density across datasets (Fig. 12).
Compared to other currently available pressure-based gait datasets, StepUP-P150 offers the highest spatial resolution at $5\times 5$ mm (Fig. 12). This represents a more than 50% increase along the axis of walking compared to the next two highest-resolution underfoot pressure databases: CAD WALK and CASIA-D. The peak pressure images in Fig. 8 illustrate the high quality of texture and shape information captured in the footstep recordings. Distinct differences in pressure distributions are visually apparent across participants, even when wearing the same shoe type, highlighting the individuality of gait patterns. Key outsole characteristics and high-pressure contact regions are also clearly visible, revealing detailed features of the participants’ personal footwear.
Time Series Representations
Biomechanical analyses also frequently involve time series signals due to the temporal nature of gait patterns, such as vertical ground reaction force (GRF) profiles, and center of pressure (COP) trajectories. GRF time series represent the forces exerted by the ground on the foot throughout the stance phase, and are computed by summing the pressures ($p$) across all pixels at each time step and converting to force in Newtons (N) using the area of each sensor ($A=2.5\times 10^{-5}\mathrm{m}^{2}$), as in (1).
$\mathrm{GRF}(t)\mathrm{\leavevmode\nobreak\ [N]}=1000\times A\mathrm{\leavevmode\nobreak\ [m^{2}]}\times\sum_{x,y}p_{xy}(t)\mathrm{\leavevmode\nobreak\ [kPa]}$ (1)
COP time series are defined in the mediolateral (ML, $x$-axis) and anteroposterior (AP, $y$-axis) directions, and are calculated as the pressure-weighted average of the foot’s coordinates at each time step during the stance, as in (2) where $s$ represents the sensor side length ($s=0.5\mathrm{\leavevmode\nobreak\ cm}$).
$\mathrm{COP_{ML}}(t)\mathrm{\leavevmode\nobreak\ [cm]}=\frac{\sum_{x,y}x\times p_{xy}(t)}{\sum_{x,y}p_{xy}(t)}\times s\mathrm{\leavevmode\nobreak\ [cm]}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{COP_{AP}}(t)\mathrm{\leavevmode\nobreak\ [cm]}=\frac{\sum_{x,y}y\times p_{xy}(t)}{\sum_{x,y}p_{xy}(t)}\times s\mathrm{\leavevmode\nobreak\ [cm]}$ (2)
The COP time series were further mean-centered to remove absolute position offsets and emphasize relative shifts in pressure distribution throughout each step.
The GRF and COP time series were used to assess data quality and compare StepUP-P150 with publicly available datasets from healthy individuals (Fig. 13). These included two datasets that used force plates (GaitRec and Gutenberg [6, 7]; Derlatka and Parfieniuk [23]) and one that used a pressure sensing walkway (CASIA-D barefoot [9]). For visualization, the StepUP-P150 signals were averaged across all footwear conditions and grouped by the four walking speeds. The resulting average GRF and COP waveforms are comparable to those from existing datasets, especially the Derlatka and Parfieniuk dataset which also included participants walking in footwear. The other two datasets primarily or exclusively included barefoot walking. Variation across walking speeds is also evident in the StepUP-P150 data, such as an increase in peak GRF during heel strike for faster walking. This is consistent with the trends observed for the CASIA-D dataset and prior biomechanics studies [24].
Spatiotemporal Gait Parameters
As shown in Fig. 14, the 120 cm $\times$ 360 cm sensing tile grid used for the StepUP-P150 dataset is substantially larger than those used in previous pressure-based gait datasets. This expanded area enables the capture of natural walking behavior across 4–6 consecutive footsteps, compared to 1–4 steps in prior datasets, and eliminates the need for subjects to target small, fixed sensing areas. As a result, spatiotemporal gait parameters such as gait speed, cadence, step length, step width, and toe-out angle can be reliably extracted from the pressure recordings. These parameters were compared to those from the CASIA-D dataset, the only other currently available public dataset with sufficient spatial resolution and sequential step count to support such analysis. Except for the toe-out angle, all features were computed from distances between 3D bounding box coordinates ($time$, $x$, $y$) of consecutive footsteps. The toe-out angle was calculated as the angle between the foot’s first principal component axis (i.e., its longest dimension) and the long axis of the tile grid, with negative values indicating inward rotation and positive values indicating outward rotation.
Notably, significant differences in step length, cadence, and toe-out angle were observed between CASIA-D and StepUP-P150 (Fig. 15, $p<0.05$ using two-sample $t$-tests). These differences likely stem from sensor size and protocol variations: (1) the wider sensing platform for StepUP-P150 may have allowed for greater toe-out angles, and (2) participants turned and sometimes paused between passes, potentially resulting in slower walking speeds that reduced step length and cadence. Indeed, participants in the StepUP-P150 dataset walked 0.21 m/s slower on average than normative speeds, stratified by sex and age group [25]. However, the StepUP-P150 dataset also exhibited higher variability across the four gait parameters compared to CASIA-D. The larger participant pool, multiple footwear conditions, and varied walking speeds in StepUP-P150 contribute to a broader and more representative sample of gait behavior.
Fig. 16 further illustrates this variability, showing average step length and step width (also called support base) for the 13 participants in the CASIA-D shod dataset and 150 participants in the StepUP-P150 dataset, each wearing two distinct pairs of personal shoes. A prior study using the CASIA-D dataset [26] found these two gait parameters to be among the most identifying features for recognizing users across changes in footwear, contributing to a classification model that achieved 90.5% identification accuracy with a single step and 99.5% with five steps. However, while these two parameters may be sufficient to distinguish the 13 CASIA-D participants with near-perfect separability, StepUP-P150 contains many individuals with overlapping values, highlighting the increased challenge and realism of person recognition in larger and more diverse populations.
@@
Footstep data are provided in NPZ (.npz) and MAT (.mat) formats, while supplementary metadata is stored in CSV files. These formats are readily compatible with standard Python and MATLAB toolboxes, facilitating easy import and analysis. The size of the Python-compatible (NPZ) dataset is approximately 50 GB, while the MATLAB-compatible (MAT) dataset is approximately 118 GB. The complete collection of UNB StepUP-P150 files can be accessed on FRDR[20](https://doi.org/10.20383/103.01285), along with scripts for importing and working with the data in both Python and MATLAB (see Code Availability). Additional Python scripts are provided to support further processing, including code for data normalization as described in the Footstep Normalization section. This script offers several normalization options (see Table 3), enabling users to apply their preferred downstream processing methods. While the dataset includes preprocessed outputs from two distinct pipelines to support quick prototyping and analysis, the script offers flexibility to explore a wide range of alternative preprocessing configurations. Finally, scripts are also provided to extract several standard gait features (e.g., peak pressure images, COP and GRF time series, and spatiotemporal gait parameters like step length and width) presented in the Technical Validation section. These are supplied to assist in benchmarking and in integrating or comparing the current dataset with other datasets.
@@
Custom scripts designed for processing and technical validation are provided in the data repository as well as on the dataset’s companion GitHub page (https://github.com/UNB-StepUP/StepUP-P150) to support ongoing improvements and optimizations. These scripts were created using MATLAB (The MathWorks, Inc., Natick, Massachusetts, United States, 2023a) and Python (Python Software Foundation, 3.11). The Python scripts require specific libraries, which are listed in the requirements.txt file. This file enables library installation through The Python Package Index (PyPI, https://pypi.org) or the Anaconda software distribution (2024.02, https://www.anaconda.com). Instructions regarding the use and execution of the custom code are provided in the accompanying README file.